# Fulde-Ferrell-Larkin-Ovchinnikov States and Topological Bogoliubov Fermi Surfaces in Altermagnets: an Analytical Study

### v1

@@ V FFLO phase boundary driven by NRSS
Figure 3: (a)-(b) Sign distribution of $w_{+}^{2}-4$ and $w_{-}^{2}-4$. (c) When $w_{+}^{2}-4>0$ and $w_{-}^{2}-4>0$, $z_{1}$ and $z_{3}$ are within the unit circle while $z_{2}$ and $z_{4}$ are outside the unit circle. At $w_{+}^{2}-4=0$, $z_{1}$ and $z_{2}$ coincide with each other at $z=1$. After that, $w_{+}^{2}-4<0$ and $w_{-}^{2}-4>0$, we have (d) where $z_{1}$ and $z_{2}$ are on the unit circle, while $z_{3}$ ($z_{4}$) is inside (outside) the unit circle. When $w_{-}^{2}-4=0$, $z_{3}$ and $z_{4}$ coincide with each other at z=1z=-1. After that, $w_{+}^{2}-4<0$ and $w_{-}^{2}-4<0$, we have (e) where all four poles are on the unit circle. In (d)-(e), small arcs (blue and red solid line) are added to the unit circle to form a closed contour (solid line) to remove the poles on the unit circle.
V.1 Contour integration of $I_{4}$
There are two integral variables $k$ and $\phi$ in $I_{4}$. The integration order can be first $k$ then $\phi$, or first $\phi$ then $k$. We find that the second choice is much more convenient than the first one and thus adopt it. Since $\bf q$} is along the $(1,1)$ direction, we have $(q_{x},q_{y})=(\frac{q}{\sqrt{2}},\frac{q}{\sqrt{2}})$ ($q>0$ for simplicity) and the denominator of the integrand of $I_{4}$ is,
@@ V.1 Contour integration of $I_{4}$
V.2 Positions of the four poles
@@
The two discriminants are calculated as $w_{\pm}^{2}-4=(1+t_{\textrm{AM}})D_{\mp}(k)/(2t_{\textrm{AM}}^{2}k^{2})$ with
@@
which are shown in Fig. 3(a)-(b). These two critical $Q$ values separate the $(k,Q)$-parameter space into different regions according to the sign of $w_{\pm}^{2}-4$. In Fig. 3(a) (Fig. 3(b)), the region marked by “+” sign means $w_{+}^{2}-4>0$ ($w_{-}^{2}-4>0$), while $w_{+}^{2}-4<0$ ($w_{-}^{2}-4<0$) in the region marked by “-” sign.
@@ V.4 Integration over $k$
Figure 4: $\frac{S(t_{\textrm{AM}},Q,0)}{\mathcal{A}}$ for $\varepsilon_{B}=0.05$ and $\mu=1.00$.
For a given $Q$, it can intersect $Q_{c1}(k)$ and $Q_{c2}(k)$ with either two or one crossings. Therefore, there are two different regimes for $Q$: (i) $Q<1/\sqrt{1-t_{\textrm{AM}}}-1/\sqrt{1+t_{\textrm{AM}}}$; and (ii) $1/\sqrt{1-t_{\textrm{AM}}}-1/\sqrt{1+t_{\textrm{AM}}}<Q<1/\sqrt{1-t_{\textrm{AM}}}+1/\sqrt{1+t_{\textrm{AM}}}$.
For $Q$ belongs to regime (i), $Q$ intersects $Q_{c1}(k)$ with two points $k_{\pm}$ whose square are given by:
@@
So $I_{4}$ is calculated as,
@@
V.5 Fixed chemical potential
With $I_{2}$, $I_{3}$, and $I_{4}$ evaluated, $S(t_{\textrm{AM}},{\bf q},0)/\mathcal{A}$ now is calculated as:
$\frac{S(t_{\textrm{AM}},Q,0)}{\mathcal{A}}=\frac{1}{4\pi}\begin{cases}\ln\sqrt{\frac{2\varepsilon_{B}}{\mu}}-\ln\left|\frac{a-b}{a+b}\right|-\ln\left|\sqrt{4-(1-t_{\textrm{AM}}^{2})Q^{2}}\right|,Q\in(0,\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}}]\\ \ln\sqrt{\frac{2\varepsilon_{B}}{\mu}}-\ln\left|\frac{a-b}{a+b}\right|+\ln\left|\frac{2t_{\textrm{AM}}}{\sqrt{C(1)}+(1-t_{\textrm{AM}}^{2})+\sqrt{2D_{+}(1)}}\right|,Q\in(\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}},\sqrt{\frac{1}{1-t_{\textrm{AM}}}}+\sqrt{\frac{1}{1+t_{\textrm{AM}}}}]\end{cases}$ (68)
Fig. 4 plot $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ for $\varepsilon_{B}=0.05$ and $\mu=1.00$. It is clear that $S(t_{\textrm{AM}},Q,0)$ increases slowly for $Q$ in regime (i) and decreases rapidly in regime (ii), therefore $Q_{c}\equiv\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}}$ is a kink for $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ which reaches the maximum value:
Figure 5: (a) Fermi surface nesting for $\mathbf{q}$ along (1, 1) direction; (b) BFS for $\mathbf{q}=(q_{c}/\sqrt{2},q_{c}/\sqrt{2})$. (c) Fermi surface nesting for $\mathbf{q}$ along (1, 0) direction. (d) BFS for $\mathbf{q}=(q_{c},0)$. Here $\Delta=0$, $\mu=1$, and $t_{AM}=0.5$ for BFS plotting.
It is well known that the N phase is stable at $S(t_{\textrm{AM}},Q,0)/\mathcal{A}<0$, therefore we see that the FFLO phase emerges at $Q=Q_{c}$ and $S(t_{\textrm{AM}},Q,0)/\mathcal{A}=0$, the simultaneous solution of which leads to the critical $t_{\textrm{AM}}$ for FFLO phase:
In Fig. 2(a), we plot the FFLO phase boundary $t_{\textrm{AM, FFLO}}$ as the red curve. When $\mu<\varepsilon_{B}$, $S(t_{\textrm{AM}},Q_{c},0)=0$ has no solution and the system remains in the N phase. For $\mu\geq\varepsilon_{B}$ the FFLO phase emerges from the N phase side of the phase diagram, in agreement with our assumption that the decrease of $t_{\textrm{AM}}$ leads to a continuous N-FFLO phase transition. From Fig. 2(a), $t_{\textrm{AM}}$ has the largest window for observing FFLO phase when $\mu=\varepsilon_{B}$. After that, the window gradually shrinks.
In Soto-Garrido et al. (Soto-Garrido2014, ), the integration over $k$ is carried out within a shell around the Fermi surface, and such a finite range of integration leads to a flat $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$. Based on such a result, Soto-Garrido et al. (Soto-Garrido2014, ) claimed the absence of FFLO phase. Our results reveal the importance of full $\mathbf{k}$-integration in restoring the FFLO phase.
V.6 Fixed total particle number
V.7 Geometric interpretation
In previous section, we have shown that, for $\Delta=0$, $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ exhibits a nonanalytic dependence on $Q$ according to the contour integration of complex analysis. The nonanalyticity occurs at $Q_{c}=\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}}$ which corresponds to $q_{c}=a-b$. As shown in Fig. 5(a), such a $q_{c}$ is the critical momentum that two of the four crossings between the two Fermi ellipses ($E_{{\bf k}\pm}=A({\bf k})\pm B({\bf k})=0$) and the black dashed circle ($A({\bf k})=0$) becomes one tangent point.
As elaborated in Appendix A, our numerical simulation verifies that Eq. (72) indeed gives a good estimate of the actual $q_{c}$ when $\mathbf{q}$ is along the $(1,0)$ direction.
VI CONCLUSIONS AND OUTLOOKS
In summary, we have analytically investigated the ground-state phase diagram of 2D spin-1/2 Fermi gases exhibiting $d$-wave altermagnetic spin splitting in the presence of $s$-wave pairing, within the framework of BdG mean-field theory. Under both fixed chemical potential and fixed total particle number conditions, we identify four distinct phases as illustrated in Fig. 2. Our results definitely confirm the existence of FFLO states under $s$-wave pairing, thereby resolving previous contradictory conclusions in the literatures. In addition, we provide a geometric interpretation of the onset $\mathbf{q}_{c}$ of FFLO states, offering intuitive insight into its origin. Quite unexpectedly, we also discover a nodal SF, characterized by the TBFS. Given the simplicity of the system - a single band, without external magnetic field, and under nodeless $s$-wave pairing - the presence of TBFS is both surprising and compelling.
Fig. 2 illustrates the fundamental relationship between anisotropic NRSS and attractive pairing. When the attractive pairing is strong, BCS state is expected with a zero-momentum and fully gapped Fermi surface. Meanwhile, for a very large anisotropic NRSS, the metallic normal phase is anticipated. Nodal SF and FFLO occur when anisotropic NRSS and attractive pairing are both at intermediate strength. All phases can be realized in either ultracold gases or electronic superconductors by choosing an appropriate $t_{\textrm{AM}}$. In the context of ultracold gases, a large $t_{\textrm{AM}}$ is required as shown in Fig. 2(a), so we can access to all four phases via tuning the two-body interactions $\varepsilon_{B}$ through Feshbach resonance Chin2010 . For electronic superconductors, a small $t_{\textrm{AM}}$ suffices, as shown in Fig. 2(b). By increasing the Fermi energy $\varepsilon_{F}$ via techniques like electrostatic gating Wu2023 , we can sequentially observe BCS, nodal SF, FFLO, and normal phases.
We conclude this section by noting several natural extensions of this work. First, incorporating additional physical ingredients - such as external magnetic fields, spin-orbit coupling, $d$-wave pairing, and finite temperature - could significantly enrich the phase diagram and reveal new emergent pairing phenomena. Second, while our analysis is based on the simplest single-plane-wave ansatz here, higher plane-wave FFLO ansatz such as the LO state with $\Delta(\mathbf{x})=\Delta\cos(\mathbf{q}\cdot\mathbf{x})$ are expected to be energetically more favorable. Finally, the topological BFS uncovered in this work may provide a fertile platform for realizing Majorana bound states, opening the door to potential applications in topological quantum computation.
Acknowledgements.
@@
Appendix A Numerical calculations of the inverse vertex function $S(t_{\textrm{AM}},\mathbf{q},\omega=0)$

### v2

@@ Appendix A Integration of $I_{4}$
Figure 5: (a)-(b) Sign distribution of $w_{+}^{2}-4$ and $w_{-}^{2}-4$ (see Eq. (61) for definition) on the ($k,Q$) domain. The two curves $Q_{c1}(k)$ and $Q_{c1}(k)$ are given by Eq. (63). (c) When $w_{+}^{2}-4>0$ and $w_{-}^{2}-4>0$, $z_{1}$ and $z_{3}$ are within the unit circle while $z_{2}$ and $z_{4}$ are outside the unit circle. At $w_{+}^{2}-4=0$, $z_{1}$ and $z_{2}$ coincide with each other at $z=1$. After that, $w_{+}^{2}-4<0$ and $w_{-}^{2}-4>0$, we have (d) where $z_{1}$ and $z_{2}$ are on the unit circle, while $z_{3}$ ($z_{4}$) is inside (outside) the unit circle. When $w_{-}^{2}-4=0$, $z_{3}$ and $z_{4}$ coincide with each other at z=1z=-1. After that, $w_{+}^{2}-4<0$ and $w_{-}^{2}-4<0$, we have (e) where all four poles are on the unit circle. In (d)-(e), small arcs (blue and red solid line) are added to the unit circle to form a closed contour (solid line) to remove the poles on the unit circle.
A.1 Contour integration of $I_{4}$
Since $\bf q$} is along the $(1,1)$ direction, we have $(q_{x},q_{y})=(\frac{q}{\sqrt{2}},\frac{q}{\sqrt{2}})$ ($q>0$ for simplicity) and the denominator of the integrand of $I_{4}$ is,
@@ A.1 Contour integration of $I_{4}$
A.2 Positions of the four poles
@@
The two discriminants are calculated as
$w_{\pm}^{2}-4=(1+t_{\textrm{AM}})D_{\mp}(k)/(2t_{\textrm{AM}}^{2}k^{2}),$ (61)
with
@@
which are shown in Fig. 5(a)-(b). These two critical $Q$ values separate the $(k,Q)$ domain into different regions according to the sign of $w_{\pm}^{2}-4$. In Fig. 5(a) (Fig. 5(b)), the region marked by ”+” sign means $w_{+}^{2}-4>0$ ($w_{-}^{2}-4>0$), while $w_{+}^{2}-4<0$ ($w_{-}^{2}-4<0$) in the region marked by ”-” sign.
@@ A.4 Integration over $k$
For a given $Q$, it can intersect $Q_{c1}(k)$ and $Q_{c2}(k)$ with either two or one crossings. Therefore, there are two domains for $Q$: (i) $Q<Q_{c}$; and (ii) $Q_{c}<Q<1/\sqrt{1-t_{\textrm{AM}}}+1/\sqrt{1+t_{\textrm{AM}}}$.
For $Q$ belongs to domain (i), $Q$ intersects $Q_{c1}(k)$ with two points $k_{\pm}$ whose square are given by:
@@
So the integration over [0, 1] now is split into three pieces $0,k_{-}]$, $k_{-},k_{+}]$, and $k_{+},1]$, where $I_{4}(k)$ takes the form $I^{(1)}_{4}(k)$, $I^{(2)}_{4}(k)$, and $I^{(3)}_{4}(k)=0$, respectively. Therefore $I_{4}$ is calculated as
@@
Appendix B Numerical calculations of the inverse vertex function $S(t_{\textrm{AM}},\mathbf{q},\omega=0)$